# Generalised Fermat equation: a survey of solved cases

### v1

@@ 1 Introduction
1.1 Generalised Fermat equation
@@
$x^{n}+y^{n}=z^{n}$ (1.1)
has no solutions in positive integers. This statement became known as Fermat’s Last Theorem (FLT), and has been one of the most famous open problems in mathematics for over $350$ years. After cancelling the common factors if needed, it suffices to prove the non-existence of positive integer solutions such that $\gcd(x,y,z)=1$. The FLT has been finally proved by Andrew Wiles [^1] With part of the proof delegated to a joint paper with Richard Taylor [138]. [144] in 1995, and this proof is widely recognised as one of the most monumental achievements of modern mathematics.
@@
where $x,y,z$ are integer variables, while positive integers $p,q,r$ and integers $a,b,c$ are parameters. If $abc=0$, then (1.2) has at most two monomials and is easy to solve [75, Section 1.6], so from now on we will assume that $abc\neq 0$. If (1.2) is solvable in integers $(x,y,z)\neq(0,0,0)$, then, under a minor condition [^2] Specifically, under the condition that at least one of the integers $p,q,r$ is coprime with another two. on $(p,q,r)$, it has infinitely many integer solutions, and all these solutions are easy to describe, see [76, Section 3.4.2] and [75, Proposition 4.53]. However, the problem of describing primitive integer solutions to (1.2), that is, ones satisfying $\gcd(x,y,z)=1$, turns out to be much more interesting and delicate.
The triple $(p,q,r)$ is called the signature of equation (1.2). If $\min\{p,q,r\}=1$ then (1.2) is trivial [75, Section 1.5.1], so we may assume that $\min\{p,q,r\}\geq 2$. The theory of primitive solutions to (1.2) crucially depends on whether the quantity $1/p+1/q+1/r$ is greater than $1$, equal to $1$, or less than $1$. In the first case, we have the following result proved by Beukers [16] in 1998.
@@
Assume that $11/p+1/q+1/r>$ 1 and $\min\{p,q,r\}\geq 2$. Then if (1.2) has at least one primitive solution with $xyz\neq 0$, then it has infinitely many such solutions. In this case, there exists a finite number of parametrised solutions of (1.2) with two parameters such that all primitive solutions can be obtained by specialisation of the parameters to integer values. Moreover, there is an algorithm that computes these parametrized solutions.
@@
In the case $11/p+1/q+1/r=$ 1, which consists of permutations of $(2,3,6)$, $(2,4,4)$ and $(3,3,3)$, equation (1.2) reduces to the problem of computing rational points on certain elliptic curves. This problem is open in general, but there are algorithms that work well for all specific equations with not-too-large coefficients $a,b,c$. For example, as early as in 1954, Selmer [121, 122] solved (1.2) with $(p,q,r)=(3,3,3)$ for all $a,b,c$ satisfying $1\leq|abc|\leq 500$. With modern computers and improved algorithms one may proceed much further if needed.
The most interesting and challenging case of equation (1.2) is $11/p+1/q+1/r<$ 1. In this case, Darmon and Granville [52] proved the following fundamental result.
@@
For any given positive integers $p,q,r,a,b,c$ satisfying $11/p+1/q+1/r<$ 1, the generalized Fermat equation (1.2) has only finitely many primitive integer solutions.
@@
$(n,n,3),n\geq 3$ [53, 114]
@@
Table 1.2: Known primitive positive integer solutions to equation (1.3) with $11/p+1/q+1/r<$ 1, up to exchange of $(x,p)$ with $(y,q)$.
$x^{p}+y^{q}=z^{r},$ (1.3)
is known as Fermat-Catalan equation, and has been particularly well-studied. A number of survey papers exist which are devoted to this equation, see, for example, [8, 90], Section 1.2 in [80], and wikipedia page https://en.wikipedia.org/wiki/Beal_conjecture. We provide a summary of solved signatures $(p,q,r)$ in Table 1.1. There is also some recent work with partial results for other signatures, see e.g [4, 7, 8, 15, 20, 21, 29, 40, 42, 43, 46, 48, 49, 62, 69, 72, 89, 94, 101, 117, 141]. Famous Fermat-Catalan conjecture predicts that (1.3) has only finitely many solutions in positive integers $(p,q,r,x,y,z)$ such that $11/p+1/q+1/r<$ 1 and $\text{gcd}(x,y,z)=1$. In fact, the only known such solutions are listed in Table 1.2, and failing to find other solutions despite extensive efforts [106, 125] suggests that this list might be complete. One may note that all exponents $(p,q,r)$ in Table 1.2 have $\min(p,q,r)=2$. In 1997, Beal [96] offered a million-dollar prize for the proof or disproof that (1.3) has no solutions in coprime positive integers $(x,y,z)$ when $\min\{p,q,r\}\geq 3$.
Comparison of Table 1.1 with Table 1.3 below shows that the only cases with $(p,q,r)$ not-all-primes that remain to be considered are $(p,q,r)=(2,5,9)$ and $(p,q,r)=(2,3,25)$. In the case when all $(p,q,r)$ are primes, Table 1.1 covers the triples $(2,3,7)$, $(5,5,7)$, $(5,5,19)$, $(7,7,5)$, $(3,3,p)$ for $p\leq 10^{9}$, $(p,p,2)$, $(p,p,3)$ and $(p,p,p)$ for all $p$, as well as $(2,3,11)$ conditional on the generalized Riemann hypothesis. Hence, the triple $(p,q,r)$ with the smallest sum $p+q+r$ for which equation (1.3) is currently open is $(p,q,r)=(2,5,7)$. The smallest triple for which Beal’s conjecture remains open is $(p,q,r)=(3,5,7)$.
While there are good surveys of equation (1.3), a similar survey for the more general equation (1.2) is not available. There are many dozens of papers studying this equation, and each of them cites just a small portion of others. As a result, some equations of the form (1.2) have been solved multiple times. For example, equation $x^{4}+3y^{4}=z^{3}$ has been solved in 1994 by Terai [139], and then again in 2019 by Söderlund [132].
The aim of this survey is to summarise the solved cases of the generalised Fermat equation (1.2) with $|abc|>1$. This will help the readers to avoid duplicate work by solving the already solved cases. To keep the survey relatively short, we will not discuss proofs. Despite this, we hope that collecting all these references in one place should make it much easier for the readers to find the proof techniques in the original papers.
In the rest of the introduction, we discuss some easy reductions, elementary special cases, and state some conjectures. Sections 2 to 5 discuss for which parameters $(a,b,c,p,q,r)$ equation (1.2) has been solved. The solved cases are then summarised in the final table in Section 6.
@@ 1.2 Easy reductions and elementary special cases
As mentioned above, we may assume that either $p,q,r$ are all primes, or $(p,q,r)$ is a permutation of one of the special triples listed in Table 1.3. In the second case, at least one of integers $p,q,r$ is composite. If, for example, $r$ is composite, and $1<d<r$ is any divisor of $r$, then change of variables $Z=z^{r/d}$ reduces (1.2) to the equation
@@
$z^{r/d}=Z=P_{i}(u,v)\quad\text{for some}\quad i\in\{1,2,\dots,k\},$ (1.5)
where $P_{1},\dots,P_{k}$ are some polynomials in two variables $u,v$ with integer coefficients. In some cases, equations (1.5) are easy to solve, which leads to the resolution of (1.2).
If $11/p+1/q+1/d=$ 1, then (1.4) reduces to computing rational points on an elliptic curve. If this curve has rank 00, then there is a finite number of such points, and they can be explicitly listed, see e.g. [75, Section 3.4.4]. Then (1.4) has a finite number of primitive solutions, and it is easy to check for which of them (if any) $z=Z^{d/r}$ is an integer.
For some of the special triples in Table 1.3 more than one of the exponents $p,q,r$ is composite, and some exponents may have more than one non-trivial divisor $d$. Hence, equation (1.2) with special $(p,q,r)$ may be reducible to several equations of the form (1.4). If any of the reduced equations has finitely many primitive solutions, and these solutions can be computed, this immediately solves (1.2). For example, this solves equation $x^{4}+2y^{4}=z^{4}$, which is the case $(a,b,c,p,q,r)=(1,2,1,4,4,4)$ of the equation (1.2), because the corresponding equation $x^{4}+2y^{4}=Z^{2}$ reduces to finding rational points on rank 00 elliptic curve $Y^{2}=X^{4}+2$. From now on, all equations (1.2) solvable by this method will be excluded from this survey.
In summary, below we will discuss the generalized Fermat’s equations (1.2), not covered by Proposition 1.5, in which $a,b,c$ are positive integers, and exponents $(p,q,r)$ are either (i) prime numbers satisfying $11/p+1/q+1/r<$ 1, or (ii) one of the special triples listed in Table 1.3 such that all the corresponding equations (1.4) have infinitely many primitive solutions. For each such equation, we will consider the following problem.
Problem 1.6.
@@
Problem 1.7.
For a given $a,b,c,p,q,r$, determine whether equation (1.2) has any non-trivial integer solutions.
Because the solutions to (1.2) with $xyz=0$ are easy to find, the negative answer to Problem 1.7 solves Problem 1.6 for the corresponding equation as well. However, if (1.2) has some easy-to-find non-trivial integer solution, this automatically solves Problem 1.7 with a “Yes” answer, without resolving Problem 1.6 for this equation.
@@ 1.3 Some conjectures and open problems
Famous and well-believed abc conjecture of Masser and Oesterlé [110] predicts that for every real number $\epsilon>0$ there exist only finitely many triples $(A,B,C)$ of coprime positive integers such that $A+B=C$ and
$C>\mathrm{rad}(ABC)^{1+\epsilon},$
where $\mathrm{rad}(n)$ denotes the product of the distinct prime factors of $n$. Darmon and Granville [52, Section 5.2] observed that the abc conjecture implies the Fermat-Catalan conjecture about the finiteness of primitive integer solutions to (1.3), and remarked that the same argument may be extended to show that it also implies the following conjectures.
@@
For any integers $a,b,c$ such that $abc\neq 0$, equation (1.2) has only finitely many solutions [^4] Solutions to (1.2) that have $\min\{x,y,z\}=1$ and differ only by the exponent(s) of $1$ are assumed to be the same, e.g. family $(x,y,z,p,q,r)=(1,1,1,p,q,r)$ for $(a,b,c)=(1,1,2)$ is considered as one solution. in integers $(x,y,z,p,q,r)$ such that $\gcd(x,y,z)=1$, $p>0$, $q>0$, $r>0$, $11/p+1/q+1/r<$ 1.
@@
For any finite set $S$ of primes, equation (1.2) has only finitely many solutions in integers $(a,b,c,x,y,z,p,q,r)$ such that $\gcd(ax,by,cz)=1$, $p>0$, $q>0$, $r>0$, $11/p+1/q+1/r<$ 1, and all the prime factors of $abc$ belongs to set $S$.
@@
Equation (1.3) corresponds to the case $a=b=c=1$ of Conjecture 1.8. The “next” case one may consider is $|abc|=2$. By rearranging (1.2) appropriately, we may assume that $a,b,c,x,y,z$ are all positive integers, and consider equation
@@
for which we are looking for solutions in positive integers $a,b,c,p,q,r,x,y,z$. Our computer search [^5] This search used the ALICE High Performance Computing facility at the University of Leicester. in the range $cz^{r}\leq 3.43\cdot 10^{22}$ returned only the solutions listed in Table 1.4.
Open Question 1.10.
Find all solutions to (1.6) in positive integers $a,b,c,p,q,r,x,y,z$. In particular, does it have any solutions except of ones listed in Table 1.4?
While we are not sure whether the solution list to (1.6) in Table 1.4 is complete, we conjecture that (1.6) has no solutions with $\min(p,q,r)\geq 3$ other than $2\cdot 5^{4}+3^{4}=11^{3}$ and $43^{4}+237^{3}=2\cdot 203^{3}$.
Another goal, suggested in [76, Section 3.4.2], is to describe all (not necessary primitive) integer solutions to (1.2), or at least to (1.3). This might be easier because for some exponents $(p,q,r)$ there exist formulas describing all integer solutions to these equations that give no hint on the structure of primitive solutions. As mentioned above, such formulas has been derived in [76] for equation (1.2) under the condition that at least one of the integers $p,q,r$ is coprime with the other two. This condition fails, if, for example, $(p,q,r)=(PQ,QR,RP)$ for some integers $P,Q,R$ greater than $1$. In this case, (1.2) reduces to $ax^{PQ}+by^{QR}=cz^{RP}$, or $a+b(y^{R}/x^{P})^{Q}=c(z^{R}/x^{Q})^{P}$. This equation can be written as
@@
where $U=z^{R}/x^{Q}$ and $V=y^{R}/x^{P}$ are rational variables. If $P+Q\geq 7$, this is an equation of genus $g\geq 2$, hence it has finitely many rational solutions by Falting’s Theorem [66], and, if we could list them all, we would be able to easily describe all possible $x,y,z$. In the case $a=b=c=1$, (1.7) reduces to $U^{P}-V^{Q}=1$. Mihăilescu [100], confirming famous conjecture of Catalan, proved that the only positive integer solutions to this equation with $P,Q\geq 2$ is $(U,V,P,Q)=(3,2,2,3)$, but solving this equation in rational $U,V$ seems to be more difficult, see [124, Theorem 12.4] and [99, Theorem 5] for a partial progress.
All conjectures and open questions mentioned in this section so far looks very difficult. Readers preferring easier-looking open problems are invited to investigate equations listed in Table 2.2, see Section 2.5.

### v2

@@ 1 Introduction
1.1 Generalized Fermat equation
@@
$x^{n}+y^{n}=z^{n}$
has no solutions in positive integers. This statement became known as Fermat’s Last Theorem (FLT), and has been one of the most famous open problems in mathematics for over $350$ years. After cancelling the common factors if needed, it suffices to prove the non-existence of positive integer solutions such that $\gcd(x,y,z)=1$. The FLT has been finally proved by Andrew Wiles [^1] With part of the proof delegated to a joint paper with Richard Taylor [143]. [149] in 1995, and this proof is widely recognized as one of the most monumental achievements of modern mathematics.
@@
where $x,y,z$ are integer variables, while positive integers $p,q,r$ and integers $a,b,c$ are parameters. If $abc=0$, then (1.1) has at most two monomials and is easy to solve, see [77, Section 1.6] and [148, Section 1.16], so from now on we will assume that $abc\neq 0$. If (1.1) is solvable in integers $(x,y,z)\neq(0,0,0)$, then, under a minor condition [^2] Specifically, under the condition that at least one of the integers $p,q,r$ is coprime with another two. on $(p,q,r)$, it has infinitely many integer solutions, and all these solutions are easy to describe, see [78, Section 3.4.2] and [77, Proposition 4.53]. However, the problem of describing primitive integer solutions to (1.1), that is, ones satisfying $\gcd(x,y,z)=1$, turns out to be much more interesting and delicate.
The triple $(p,q,r)$ is called the signature of Eq. (1.1). If $\min\{p,q,r\}=1$ then (1.1) is trivial, see [77, Section 1.5.1] and [148, Section 1.12], so we may assume that $\min\{p,q,r\}\geq 2$. The theory of primitive solutions to (1.1) crucially depends on whether the quantity $1/p+1/q+1/r$ is greater than $1$, equal to $1$, or less than $1$. In the first case, we have the following result proved by Beukers [18] in 1998.
@@
Assume that $11/p+1/q+1/r>$ 1 and $\min\{p,q,r\}\geq 2$. Then if (1.1) has at least one primitive solution with $xyz\neq 0$, then it has infinitely many such solutions. In this case, there exists a finite number of parametrized solutions of (1.1) with two parameters such that all primitive solutions can be obtained by specialization of the parameters to integer values. Moreover, there is an algorithm that computes these parametrized solutions.
@@
In the case $11/p+1/q+1/r=$ 1, which consists of permutations of $(2,3,6)$, $(2,4,4)$ and $(3,3,3)$, Eq. (1.1) reduces to the problem of computing rational points on certain elliptic curves. This problem is open in general, but there are algorithms that work well for all specific equations with not-too-large coefficients $a,b,c$. For example, as early as in 1954, Selmer [126, 127] solved (1.1) with $(p,q,r)=(3,3,3)$ for all $a,b,c$ satisfying $1\leq|abc|\leq 500$. With modern computers and improved algorithms one may proceed much further if needed.
The most interesting and challenging case of Eq. (1.1) is $11/p+1/q+1/r<$ 1. In this case, Darmon and Granville [54] proved the following fundamental result.
@@
For any given positive integers $p,q,r,a,b,c$ satisfying $11/p+1/q+1/r<$ 1, the generalized Fermat Eq. (1.1) has only finitely many primitive integer solutions.
@@
$(n,n,3),n\geq 3$ [55, 117] [^4] It was noted on Mathoverflow (https://mathoverflow.net/questions/488724/) that [55] does not provide details for the case $(3,4,4)$, this case is proven in [47, Proposition 14.6.6].
@@
$x^{p}+y^{q}=z^{r},$ (1.2)
is known as the Fermat–Catalan equation, and has been particularly well-studied. A number of survey papers exist which are devoted to this equation, see, for example, [9, 92], Section 1.2 in [82], and wikipedia page https://en.wikipedia.org/wiki/Beal_conjecture. We provide a summary [^5] Recently, Bartolomé and Mihăilescu [3] claimed to solve (1.2) for all signatures $(p,p,r)$ such that $p,r$ are primes, $p\geq 5$, and $r>3\sqrt{p\log_{2}p}$. To the best of our knowledge, this paper has not been peer reviewed yet. of solved signatures $(p,q,r)$ in Table 1.1. There is also some recent work with partial results for other signatures, see e.g [5, 8, 9, 17, 21, 22, 30, 41, 43, 44, 48, 49, 50, 51, 60, 71, 74, 91, 96, 103, 121, 146]. Famous Fermat–Catalan conjecture predicts that (1.2) has only finitely many solutions in positive integers $(p,q,r,x,y,z)$ such that $11/p+1/q+1/r<$ 1 and $\text{gcd}(x,y,z)=1$. In fact, the only known such solutions are listed in Table 1.2, and failing to find other solutions despite extensive efforts [109, 130] suggests that this list might be complete. One may note that all exponents $(p,q,r)$ in Table 1.2 have $\min(p,q,r)=2$. In 1997, Beal [98] offered a million-dollar prize for the proof or disproof that (1.2) has no solutions in coprime positive integers $(x,y,z)$ when $\min\{p,q,r\}\geq 3$.
Table 1.2: Known primitive positive integer solutions to (1.2) with $11/p+1/q+1/r<$ 1, up to exchange of $(x,p)$ with $(y,q)$.
@@
Comparison of Table 1.1 with Table 1.3 below shows that the only cases of Eq. (1.2) with $(p,q,r)$ not-all-primes that remain to be considered are $(p,q,r)=(2,5,9)$ and $(p,q,r)=(2,3,25)$. In the case when all $(p,q,r)$ are primes, Table 1.1 covers the triples $(2,3,7)$, $(5,5,7)$, $(5,5,19)$, $(7,7,5)$, $(3,3,p)$ for $p\leq 10^{9}$, $(p,p,2)$, $(p,p,3)$ and $(p,p,p)$ for all $p$, as well as $(2,3,11)$ conditional on the generalized Riemann hypothesis. Hence, the triple $(p,q,r)$ with the smallest sum $p+q+r$ for which Eq. (1.2) is currently open is $(p,q,r)=(2,5,7)$. The smallest triple for which Beal’s conjecture remains open is $(p,q,r)=(3,5,7)$.
Recently, Sikora [130] reported that Eq. (1.2) has no solutions in coprime positive integers other than ones listed in Table 1.2 in the range $z^{r}<2^{71}$. Restricting the search to only exponents not covered in Table 1.1 allowed us to significantly speed up the search and reach the bound $2^{100}$.
Proposition 1.3.
The only solutions to (1.2) in coprime positive integers such that $z^{r}\leq 2^{100}$ are those listed in Table 1.2.
While there are good surveys of Eq. (1.2), a similar survey for the more general Eq. (1.1) is not available. There are many dozens of papers studying this equation, and each of them cites just a small portion of others. As a result, some equations of the form (1.1) have been solved multiple times. For example, equation $x^{4}+3y^{4}=z^{3}$ has been solved in 1994 by Terai [144], and then again in 2019 by Söderlund [137].
The aim of this survey is to summarize the solved cases of the generalized Fermat Eq. (1.1) with $|abc|>1$. This will help the readers to avoid duplicate work by solving the already solved cases. To keep the survey relatively short, we will not discuss proofs. Despite this, we hope that collecting all these references in one place should make it much easier for the readers to find the proof techniques in the original papers.
In the rest of the introduction, we discuss some easy reductions, elementary special cases, and state some conjectures. Sections 2 to 5 discuss for which parameters $(a,b,c,p,q,r)$ Eq. (1.1) has been solved. The solved cases are then summarized in the final table (Table 6.1) in Section 6.
@@ 1.2 Easy reductions and elementary special cases
As mentioned above, we may assume that either $p,q,r$ are all primes, or $(p,q,r)$ is a permutation of one of the special triples listed in Table 1.3. In the second case, at least one of the integers $p,q,r$ is composite. If, for example, $r$ is composite, and $1<d<r$ is any divisor of $r$, then change of variables $Z=z^{r/d}$ reduces (1.1) to the equation
@@
$z^{r/d}=Z=P_{i}(u,v)\text{ for some }i\in\{1,2,\dots,k\},$ (1.4)
where $P_{1},\dots,P_{k}$ are some polynomials in two variables $u,v$ with integer coefficients. In some cases, Eqs. (1.4) are easy to solve, which leads to the resolution of (1.1).
If $11/p+1/q+1/d=$ 1, then (1.3) reduces to computing rational points on an elliptic curve. If this curve has rank 00, then there is a finite number of such points, and they can be explicitly listed, see e.g. [77, Section 3.4.4] and [148, Section 3.21]. Then (1.3) has a finite number of primitive solutions, and it is easy to check for which of them (if any) $z=Z^{d/r}$ is an integer.
For some of the special triples in Table 1.3 more than one of the exponents $p,q,r$ is composite, and some exponents may have more than one non-trivial divisor $d$. Hence, Eq. (1.1) with special $(p,q,r)$ may be reducible to several equations of the form (1.3). If any of the reduced equations has finitely many primitive solutions, and these solutions can be computed, this immediately solves (1.1). For example, this solves equation $x^{4}+2y^{4}=z^{4}$, which is the case $(a,b,c,p,q,r)=(1,2,1,4,4,4)$ of Eq. (1.1), because the corresponding equation $x^{4}+2y^{4}=Z^{2}$ reduces to finding rational points on rank 00 elliptic curve $Y^{2}=X^{4}+2$, where $Y=\frac{Z}{y^{2}}$, $X=\frac{x}{y}$. From now on, all Eqs. (1.1) solvable by this method will be excluded from this survey.
In summary, below we will discuss the generalized Fermat’s Eqs. (1.1), not covered by Proposition 1.6, in which $a,b,c$ are positive integers with $abc>1$, and exponents $(p,q,r)$ are either (i) prime numbers satisfying $11/p+1/q+1/r<$ 1, or (ii) one of the special triples listed in Table 1.3 such that all the corresponding Eqs. (1.3) have infinitely many primitive solutions. For each such equation, we will consider the following problem.
Problem 1.7.
@@
Problem 1.8.
For a given $a,b,c,p,q,r$, determine whether Eq. (1.1) has any non-trivial primitive integer solutions.
Because the solutions to (1.1) with $xyz=0$ are easy to find, the negative answer to Problem 1.8 solves Problem 1.7 for the corresponding equation as well. However, if (1.1) has some easy-to-find non-trivial primitive integer solution, this automatically solves Problem 1.8 with a “Yes” answer, without resolving Problem 1.7 for this equation.
@@ 1.3 Some conjectures and open problems
Famous and well-believed abc conjecture of Masser and Oesterlé [113] predicts that for every real number $\epsilon>0$ there exist only finitely many triples $(A,B,C)$ of coprime positive integers such that
$A+B=C\quad\text{and}\quad C>\mathrm{rad}(ABC)^{1+\epsilon},$ (1.5)
where $\mathrm{rad}(n)$ denotes the product of the distinct prime factors of $n$. Darmon and Granville [54, Section 5.2] observed that the abc conjecture implies the Fermat–Catalan conjecture about the finiteness of primitive integer solutions to (1.2), and remarked that the same argument may be extended to show that it also implies the following conjectures.
@@
For any integers $a,b,c$ such that $abc\neq 0$, Eq. (1.1) has only finitely many solutions [^6] Solutions to (1.1) that have $\min\{x,y,z\}=1$ and differ only by the exponent(s) of $1$ are assumed to be the same, e.g. family $(x,y,z,p,q,r)=(1,1,1,p,q,r)$ for $(a,b,c)=(1,1,2)$ is considered as one solution. in integers $(x,y,z,p,q,r)$ such that $\gcd(x,y,z)=1$, $p>0$, $q>0$, $r>0$, $11/p+1/q+1/r<$ 1.
@@
For any finite set $S$ of primes, Eq. (1.1) has only finitely many solutions in integers
$(a,b,c,x,y,z,p,q,r)$ such that $\gcd(ax,by,cz)$ $=1$, $p>0$, $q>0$, $r>0$, $11/p+1/q+1/r<$ 1, and all the prime factors of $abc$ belong to set $S$.
@@
Eq. (1.2) corresponds to the case $a=b=c=1$ of Conjecture 1.9. The “next” case one may consider is $|abc|=2$. By rearranging (1.1) appropriately, we may assume that $a,b,c,x,y,z$ are all positive integers, and consider equation
@@
for which we are looking for solutions in positive integers $a,b,c,p,q,r,x,y,z$. Our computer search [^7] This search used the ALICE High Performance Computing facility at the University of Leicester. in the range $cz^{r}\leq 2^{80}$ returned only the solutions listed in Table 1.4.
Open Question 1.11.
Find all solutions to (1.6) in positive integers $a$, $b$, $c$, $p$, $q$, $r$, $x$, $y$, $z$. In particular, does it have any solutions except of the ones listed in Table 1.4?
While we are not sure whether the solution list to (1.6) in Table 1.4 is complete, we conjecture that (1.6) has no solutions with $\min(p,q,r)\geq 3$ other than $1^{p}+1^{q}=2\cdot 1^{r}$, $2\cdot 5^{4}+3^{4}=11^{3}$ and $43^{4}+237^{3}=2\cdot 203^{3}$.
Another goal, suggested in [78, Section 3.4.2], is to describe all (not necessary primitive) integer solutions to (1.1), or at least to (1.2). This might be easier because for some exponents $(p,q,r)$ there exist formulas describing all integer solutions to these equations that give no hint on the structure of primitive solutions. As mentioned above, such formulas have been derived in [78] for Eq. (1.1) under the condition that at least one of the integers $p,q,r$ is coprime with the other two. This condition fails, if, for example, $(p,q,r)=(PQ,QR,RP)$ for some integers $P,Q,R$ greater than $1$. In this case, (1.1) reduces to $ax^{PQ}+by^{QR}=cz^{RP}$, or $a+b(y^{R}/x^{P})^{Q}=c(z^{R}/x^{Q})^{P}$. This equation can be written as
@@
where $U=z^{R}/x^{Q}$ and $V=y^{R}/x^{P}$ are rational variables. If $P+Q\geq 7$, this is an equation of genus $g\geq 2$, hence it has finitely many rational solutions by Falting’s Theorem [68], and, if we could list them all, we would be able to easily describe all possible $x,y,z$. In the case $a=b=c=1$, (1.7) reduces to $U^{P}-V^{Q}=1$. Mihăilescu [101], confirming famous conjecture of Catalan, proved that the only positive integer solutions to this equation with $P,Q\geq 2$ is $(U,V,P,Q)=(3,2,2,3)$, but solving this equation in rational $U,V$ seems to be more difficult, see [129, Theorem 12.4] and [102, Theorem 5] for a partial progress.
All conjectures and open questions mentioned in this section look very difficult. [^8] In 2022, Mochizuki et al. [104] claimed to prove the effective abc conjecture, stating that (1.5) has only finitely many solutions in coprime positive integers $(A,B,C)$, and, moreover, there is an explicitly computable upper bound for the size of any such solution. If true, this would imply the truth of Conjectures 1.9 and 1.10, the full resolution of (1.2) for all but finitely many explicit signatures [155], reduce Open Question 1.11 to finite computation, and solve many other cases of (1.1). However, Scholze and Stix [125] have found a serious flaw in the argument. Readers preferring easier-looking open problems are invited to investigate equations listed in Table 2.2, see Section 2.5.